# EgoMI: Learning Active Vision and Whole-Body Manipulation from Egocentric Human Demonstrations

### v1

@@ Abstract
Imitation learning from human demonstrations offers a promising approach for robot skill acquisition, but egocentric human data introduces fundamental challenges due to the embodiment gap. During manipulation, humans actively coordinate head and hand movements, continuously reposition their viewpoint and use pre-action visual fixation search strategies to locate relevant objects. These behaviors create dynamic, task-driven head motions that static robot sensing systems cannot replicate, leading to a significant distribution shift that degrades policy performance. We present EgoMI (Egocentric Manipulation Interface), a framework that captures synchronized end-effector and active head trajectories during manipulation tasks, resulting in data that can be retargeted to compatible semi-humanoid robot embodiments. To handle rapid and wide-spanning head viewpoint changes, we introduce a memory-augmented policy that selectively incorporates historical observations. We evaluate our approach on a bimanual robot equipped with an actuated camera head and find that policies with explicit head-motion modeling consistently outperform baseline methods. Results suggest that coordinated hand–eye learning with EgoMI effectively bridges the human-robot embodiment gap for robust imitation learning on semi-humanoid embodiments. Project page: https://egocentric-manipulation-interface.github.io
@@ I INTRODUCTION
Learning from large-scale human demonstrations represents a powerful pathway toward scalable robot skill acquisition. Recent advances in imitation learning have shown promising results in training robots directly from human collected data. However, many critical challenges emerge from the embodiment gap between humans and robots. During manipulation tasks, humans naturally integrate head movements with hand actions, dynamically adjusting their viewpoint to maintain visual contact with objects, anticipate occlusions, and retain spatial information. This active perception strategy is fundamental to how we understand and interact with our environment.
@@
In contrast, many robotic systems rely on fixed external or wrist-camera views that cannot replicate this coordinated visual behavior. When learning from egocentric human demonstrations, this embodiment gap creates severe distribution shifts, as the dynamic viewpoint changes present in human data cannot be reproduced by static robot perception systems. Various methods attempt to minimize the embodiment gap through strategies such as restricting to wrist mounted cameras [1, 2, 3, 4] or projecting top camera views into a coordinate-invariant representation [5, 6]. However, for more complex tasks that require searching or looking with the head, the learned policy is unable to reproduce the demonstrated behavior. The inability for robot policies to remember what it has seen in the past further adds to this gap. Together, these mismatches create substantial embodiment gaps that hinder policy learning and degrade imitation performance.
@@
To address this challenge, we introduce Egocentric Manipulation Interface (EgoMI), as illustrated in LABEL:fig:teaser, a framework that captures the full degrees of freedom of human perceptual movements with minimal embodiment gap. EgoMI simultaneously records both end-effector and head movements during human demonstrations. Proprioception and camera data streams for both the wrist and head result in more complete conditioning information for the downstream policy. We can then retarget a policy’s prediction of the whole-body motion data onto a robot. We use a modified Rainbow-RBY1 (wheeled, semi humanoid robot), equipped with a 6-DoF arm that functions as a neck to enable faithful retargeting of human motions to the robot.
@@
Additionally, we present Spatial-Aware Robust Keyframe Selection (SPARKS), a lightweight algorithm that selects a compact set of past head keyframe images to mitigate the loss of context caused by rapid head viewpoint shifts, while avoiding out-of-distribution failures. SPARKS is a simple yet effective mechanism that emphasizes keyframes rich in spatial information, and scores past frames using viewpoint novelty, temporal recency, and motion smoothness as a training-free proxy for visual informativeness. By embedding spatial memory into robot policies, SPARKS enables more stable long-horizon reasoning and resilience to viewpoint shifts.
@@
Through real world robotics experiments, we show that incorporating head observations and movement from the egocentric demonstration is crucial during both the data collection stage and policy deployment stage. Policies trained without the head trajectory data or memory consistently fail due to a lack of critical context. Remarkably, our framework achieves zero-robot-data transfer without relying on any augmentations or viewpoint re-rendering, demonstrating that capturing egocentric head and hand motions and observations is sufficient to bridge the embodiment gap. In summary, we:
@@ II-B Imitation learning approaches from human data
Recent advances in large-scale imitation learning have resulted in robot foundation model approaches that integrate pre-trained vision and language representations with behavior cloning to scale imitation learning across numerous tasks  [19, 20, 21, 22, 23]. Follow-on systems such as RoboFlamingo [24], Octo [25], OpenVLA [26] $\pi$[27] [28], $\pi$0.5[29], and Seed GR-3 [30] confirm that combining demonstration data with foundation models yields strong zero-shot generalization and versatile open-world behavior.
@@ II-C Policy learning with Active Vision
ViA [35] uses a 6-DoF head-mounted camera to imitate human head motion, reducing occlusions and selecting informative viewpoints, but leverages a robot embodiment specific teleoperation data collection system. EyeRobot [36] equips a robotic system with a 2DoF “eyeball” camera and learns gaze behavior through a combined BC-RL (behavior cloning and reinforcement learning) objective, while manipulation is learned through behavior cloning. Look, Focus, Act [37] incorporates human gaze data into a foveated transformer, showing how explicit fixation signals improve policy robustness. In contrast, EgoMI directly captures human head motion for natural active vision and combines it with SPARKS, a lightweight memory mechanism to learn policies in a robot agnostic manner while still maintaining a minimal embodiment gap.
@@ III-B Operator Gaze and Active Vision Data
Centering provides significant benefits for downstream policies by driving task-relevant visual features into the middle of the observation space and creating an object-centric representation that couples perception and action. In contrast, fixed external-cameras or un-actuated head-cameras scatter task-relevant visual features across the image plane, forcing models to rely on weaker positional encodings. Qualitatively, we observe that policies trained without the reticle often fail completely, likely due to free gaze variability and weak head–gaze correlation, which caused trajectories and observations to drift out of the demonstration distribution at deployment.
@@ IV-A Policy Interface: 29D Action and State Representation
where $r_{\cdot}^{(\cdot)}\in\mathbb{R}^{6}$ is a 6D rotation representation (first two columns of the rotation matrix), $p_{\cdot}^{(\cdot)}\in\mathbb{R}^{3}$ is position, and $g^{(\cdot)}\in\mathbb{R}$ is the continuous gripper signal. For actions, $g^{L},g^{R}$ come from operator intent (VR triggers); for state, they come from measured hardware closure/aperture.
@@
We then parameterize $({}^{R}\!T_{L},{}^{R}\!T_{H})$ by $(r^{L\!:\!R}_{6},p^{L\!:\!R}_{3})$ and $(r^{H\!:\!R}_{6},p^{H\!:\!R}_{3})$ using the same 6D rotation + 3D position convention:
@@
IV-B SPARKS: Spatial-Aware Robust Keyframe Selection
@@ IV-C Policy Training
Starting with a capable pre-trained foundation model is crucial for achieving strong final performance. We initialize our approach from pre-trained $\pi$0 model weights [27]. However, $\pi$0 was originally trained only on absolute robot joint positions, which introduces a mismatch with our target representation. To bridge this gap and map $\pi$0’s output space into the relative Cartesian space, we adopt a two-stage finetuning process:
@@ IV-D Policy Deployment
The model runs at approximately 40 ms end-to-end latency per forward pass on an RTX 5090 GPU, and we set the action control frequency to 25 Hz. The policy outputs a horizon of 40 future end-effector poses and gripper commands. We adopt temporal chunk ensembling [41] as our default control strategy. Safety and embodiment-specific kinematics are handled below the policy layer: actions are saturated (e.g., clamping gripper commands) and whole-body joint commands are computed through a differentiable inverse kinematics (IK) backend (Pyroki) [42] before being sent to the robot hardware for joint-control. The solver retargets the policy’s 29D action vector comprising of SE(3) pose targets for the left hand, right hand, and head (and 2 gripper actions) into robot joint configurations. The solver minimizes a weighted cost over end-effector errors, and posture regularization, subject to joint limits, enabling embodiment generality: the same 29D action representation can be transferred to any robot with dual manipulators and an actuated head without retraining. Unlike analytical IK, which fails on unreachable targets, the differentiable solver provides best-effort, cost-minimized solutions, allowing unconstrained VR-collected demonstrations to be mapped gracefully onto robots with physical kinematic limits.

### v2

@@ Abstract
Imitation learning from human demonstrations offers a promising approach for robot skill acquisition, but egocentric human data introduces fundamental challenges due to the embodiment gap. During manipulation, humans actively coordinate head and hand movements, continuously reposition their viewpoint and use pre-action visual search strategies to locate task-relevant objects. These behaviors create dynamic, task-driven head motions that static robot sensing systems cannot replicate, leading to a significant distribution shift that degrades policy performance. We present EgoMI (Egocentric Manipulation Interface), a framework that captures synchronized end-effector and active head trajectories during manipulation tasks, resulting in data that can be retargeted to compatible semi-humanoid robot embodiments. To handle rapid and wide-spanning head viewpoint changes, we introduce a memory-augmented policy that selectively incorporates context from historical observations. We evaluate our approach on a bimanual robot equipped with an actuated camera head and find that policies with explicit head-motion modeling consistently outperform baseline methods. Results suggest that coordinated hand–eye learning with EgoMI effectively bridges the human-robot embodiment gap for robust imitation learning on semi-humanoid embodiments. Project page: https://egocentric-manipulation-interface.github.io
@@ I INTRODUCTION
Learning from large-scale human demonstrations represents a powerful pathway toward scalable robot skill acquisition. Recent advances in imitation learning have shown promising results in training robots directly from human collected data. However, a significant barrier persists in the form of the embodiment gap, the fundamental mismatch between human demonstrators and robotic platforms. During manipulation tasks, humans actively integrate head and eye movements with hand actions, dynamically adjusting their viewpoint to maintain visual contact with task-relevant objects and resolve occlusions. This active perception strategy is fundamental to how we understand and interact with our environment.
@@
In contrast, many contemporary robotic systems rely on static-external camera configurations that cannot replicate this coordinated visual behavior. This creates a severe distribution shift when learning from egocentric data, as the task-driven viewpoint changes inherent in human demonstrations cannot be reproduced by fixed sensing systems. Various methods attempt to minimize the embodiment gap through strategies such as restricting to wrist mounted cameras [35, 6, 31, 40] or projecting top camera views into a coordinate-invariant representation [5, 36]. However, for more complex tasks that require searching or looking with the head, the learned policy is unable to reproduce the demonstrated behavior. Furthermore, the inability of standard policies to maintain a persistent spatial memory of past observations exacerbates this gap, leading to context loss during rapid head movements.
@@
To address this challenge, we introduce the Egocentric Manipulation Interface (EgoMI), as illustrated in LABEL:fig:teaser, a framework that captures the full degrees of freedom of human perceptual movements with minimal embodiment gap. EgoMI simultaneously records both end-effector and head movements during human demonstrations. Proprioception and camera data streams for both the wrist and head result in more complete conditioning information for the downstream policy. We can then retarget a policy’s prediction of the whole-body motion data onto a robot. We use a modified Rainbow-RBY1 (wheeled, semi humanoid robot), equipped with a 6-DoF arm that functions as a neck to enable faithful retargeting of human head and end-effector motions to the robot.
@@
A primary challenge of utilizing an actuated head is the potential for context loss due to large viewpoint shifts. To address this, we present Spatial-Aware Robust Keyframe Selection (SPARKS), a lightweight algorithm that selects a compact set of past head keyframe images to mitigate the loss of context caused by rapid head viewpoint shifts, while avoiding out-of-distribution failures. SPARKS is a simple yet effective mechanism that emphasizes keyframes rich in spatial information, and scores past frames using viewpoint novelty, temporal recency, and motion smoothness as a training-free proxy for visual informativeness. By embedding spatial memory into robot policies, SPARKS enables more stable long-horizon reasoning and resilience to viewpoint shifts.
@@
Real world robot experiments show that incorporating head observations and movement from the egocentric demonstration is crucial during both the data collection stage and policy deployment stage. Policies trained without the head trajectory data or memory consistently fail due to a lack of critical context. Remarkably, our framework achieves zero-robot-data transfer without relying on any augmentations, in-painting, or viewpoint re-rendering, demonstrating that capturing egocentric head and hand motions and observations is sufficient to bridge the embodiment gap. In summary, we:
@@ II-B Imitation learning approaches from human data
Recent advances in large-scale imitation learning have resulted in robot foundation model approaches that integrate pre-trained vision and language representations with behavior cloning to scale imitation learning across numerous tasks  [27, 17, 3, 9, 26]. Follow-on systems such as RoboFlamingo [23], Octo [28], OpenVLA [21] $\pi_{0}$ [2] [8], $\pi_{0.5}$[16], and Seed GR-3 [4] confirm that combining demonstration data with foundation models yields strong zero-shot generalization and versatile open-world behavior.
@@ II-C Policy learning with Active Vision
ViA [38] uses a 6-DoF head-mounted camera to imitate human head motion, reducing occlusions and selecting informative viewpoints, but leverages a robot embodiment specific teleoperation data collection system. EyeRobot [19] equips a robotic system with a 2DoF “eyeball” camera and learns gaze behavior through a combined BC-RL (behavior cloning and reinforcement learning) objective, while manipulation is learned through behavior cloning. Look, Focus, Act [7] incorporates human gaze data into a foveated vision encoder, showing how explicit fixation signals improve policy robustness. In contrast, EgoMI directly captures human head motion for natural active vision and combines it with SPARKS, a lightweight memory mechanism to learn policies in a robot agnostic manner while still maintaining a minimal embodiment gap.
@@ III-B Operator Gaze and Active Vision Data
Centering provides significant benefits for downstream policies by driving task-relevant visual features into the middle of the observation space and creating an object-centric representation that couples perception and action. In contrast, fixed external-cameras or unactuated head-cameras scatter task-relevant visual features across the image plane, forcing models to rely on weaker positional encodings. Qualitatively, we observe that policies trained without the reticle often fail completely, likely due to free gaze variability and weak head–gaze correlation, which caused trajectories and observations to drift out of the demonstration distribution at deployment.
@@ IV-A Policy Interface: 29D Action and State Representation
where $r_{\cdot}^{(\cdot)}\in\mathbb{R}^{6}$ is a 6D rotation vector (first two columns or rows of the $\mathbb{R}^{3\times 3}$ rotation matrix concatenated), $p_{\cdot}^{(\cdot)}\in\mathbb{R}^{3}$ is position, and $g^{(\cdot)}\in\mathbb{R}$ is the continuous gripper signal. For actions, $g^{L},g^{R}$ come from operator intent (VR triggers); for state, they come from measured hardware closure/aperture.
@@
We then parameterize $({}^{R}\!T_{L},{}^{R}\!T_{H})$ by $(r^{L\!:\!R}_{6},p^{L\!:\!R}_{3})$ and $(r^{H\!:\!R}_{6},p^{H\!:\!R}_{3})$ using the same 6D rotation + 3D position ordering convention:
@@
$SO(3)$ Representation
We choose this 6D vector representation for $SO(3)$ over other discontinuous representations such as Euler angles or double cover representations like quaternions, critical for the stability of neural network gradient-based optimization [43].
IV-B Spatial-Aware Robust Keyframe Selection (SPARKS)
@@ IV-C Policy Training
Starting with a capable pre-trained foundation model is crucial for achieving strong final performance. We initialize our approach from pre-trained $\pi_{0}$ model weights [2]. However, $\pi_{0}$ was originally trained only on absolute robot joint positions, which introduces a mismatch with our target representation. To bridge this gap and map $\pi_{0}$’s output space into the relative Cartesian space, we adopt a two-stage finetuning process:
@@ IV-D Policy Deployment
The action control frequency is set to 25Hz while the policy model asynchronously inferences at approximately 40 ms latency on an RTX 5090 GPU. The policy outputs a horizon of 40 future end-effector and head poses and gripper commands. We adopt temporal chunk ensembling [42] as our default control strategy.
To map these targets (left hand, right hand, and head) into the robot’s specific joint configurations, we employ a differentiable inverse kinematics (IK) solver (Pyroki) [20]. Unlike analytical IK, which fails or returns nulls on unreachable poses, a differentiable approach treats IK as a weighted cost-minimization problem. By optimizing for end-effector pose error alongside posture regularization, the system achieves ”graceful degradation” reaching as close as physically possible to the demonstrated pose rather than experiencing execution errors. This ensures robust transfer of unconstrained human demonstrations to robot embodiments with varying kinematic limits without requiring manual trajectory filtering or retraining.
We maintain hardware modularity by using the Viser [39] toolkit for high-level scene-graph management. Rather than authoring a monolithic URDF, we programmatically align the mounting interfaces of the Rainbow RBY1 and I2RT YAM actuated neck within a global coordinate frame. During deployment, two independent Pyroki processes resolve the torso/arm and head trajectories in parallel. This decoupled architecture ensures the 29D action representation remains platform-agnostic, facilitating zero-shot transfer to heterogeneous hardware configurations without retraining.